# Adaptive Draft-Verification for Efficient Large Language Model Decoding

### v1

@@ 3 Theoretical Insight: Why ADED uses MCTS
where $P$ denotes the conditional probability of LLM, $w$ denotes the tokens, and $\theta$ denotes the model parameters. Regularization techniques, such as KL divergence, are often incorporated to prevent overfitting and ensure generalization:
$\mathcal{L}(\theta)=-\sum_{t=1}^{T}\log P(w_{t}\mid w_{1:t-1};\theta)+\lambda\text{KL}(P_{\text{model}},P_{\text{data}}).\vspace{-2mm}$ (7)
Comparison between MCTS & LLM Inference: As shown in Eq (5) and Eq. (7), both MCTS and LLM Inference can be viewed as regularized optimization problems for selecting the distribution of the next tokens. On the one hand, the Q-function in MCTS for ADED can be viewed as an approximation to the log-likelihood of LLM
$Q(x,a)=-\sum_{t=1}^{T}\log\hat{P}(w_{t}\mid w_{t-1},w_{t-2};\theta)\approx P(w_{t}\mid w_{t-1},w_{t-2};\theta)\approx-\sum_{t=1}^{T}\log P(w_{t}\mid w_{1:t-1};\theta)\vspace{-2mm}$ (8)
where $\hat{P}$ and $P$ are the conditional probability distribution from tri-gram-matrix-based LLM representative and LLM, respectively. On the other hand, both MCTS and LLM Inference improve the optimization procedure by employing regularization techniques. Through a comparative analysis, we show the similarities between MCTS and LLM Inference in terms of optimization and regularization, and highlight our rationale for choosing MCTS for the ADED framework.
4 Experiments
4.1 Experimental Setup
Models and Datasets  To evaluate the efficacy of ADED for the inference of large-language models, we execute a series of experiments employing five distinct models in four datasets. We test our algorithm on three Vicuna models Chiang et al. (2023) (7B, 13B, 33B) and two LLaMA2-chat models Touvron et al. (2023b) (7B, 13B) to evaluate its acceleration capabilities across different sizes and types of models. Our assessment incorporated the HumanEval Chen et al. (2021), MT-Bench Zheng et al. (2023) and Alpaca Taori et al. (2023) dataset to ascertain general natural language understanding and generation competencies. These datasets were meticulously chosen to guarantee a comprehensive analysis of our acceleration techniques across various tasks.
Corpus  We constructe two corpus. The first one is built using a portion of the Python pre-training code from The Stack Kocetkov et al. (2022), comprising approximately 2.7M Python code samples with a resulting size of 1007MB. The second corpus is constructed using data derived from UltraChat Ding et al. (2023), consisting of around 774K ChatGPT conversations, which produces a corpus with a size of 574MB.
Metrics  To assess the performance of our acceleration algorithms on large language models, we utilize two main metrics: speedup ratio and average acceptance length. The speedup ratio, calculated as the ratio of the time required by the baseline models to complete inference tasks without acceleration to the time required by our ADED, effectively measures the efficiency gains introduced by our algorithm. The second metric, average acceptance length, measures the average number of tokens accepted per forward pass by the target large language models, excluding any overhead of retrieving and constructing draft tokens, indicating the maximum possible acceleration.
Baselines  In this study, we investigate various foundational approaches to improve the decoding speed of large language models. We examine Lookahead Decoding Fu et al. (2024a), an innovative, precise, parallel decoding algorithm that significantly cuts down latency without relying on draft models. We also assess REST He et al. (2023) (Retrieval-Based Speculative Decoding), which adopts a retrieval-based strategy to create draft tokens, in contrast to conventional speculative decoding methods that rely on a draft model. Collectively, these baseline methods provide a solid framework for evaluating the efficiency of our proposed acceleration techniques in the LLM decoding process. All experiments are conducted on an NVIDIA A6000, except for the 33B model, which utilizes an NVIDIA A100. Where not mentioned, the experiments default to Greedy sampling.
4.2 Main Experimental Results
In our experiment, we investigate the efficacy of diverse methodologies applied to multiple models, utilizing three distinct datasets: MT-Bench, Human-Eval, and Alpaca. We focus on metrics such as Accepted Length, Latency, and Speed Up to evaluate the efficiency and responsiveness.
@@ 4.2 Main Experimental Results
Table 1: Latency Comparison of ADED and Baselines. ADED gets the lowest latency in almost all the test cases, having a higher efficiency.
MT-Bench Latency Alpaca Latency
Model REST REST Single Lookahead ADED REST REST Single Thread Lookahead ADED
vicuna-7B 16.05 16.31 20.53 12.95 13.21 13.51 20.86 13.18
vicuna-13B 25.43 25.99 33.55 22.94 22.27 22.65 34.62 22.80
vicuna-33B 32.90 33.26 42.13 29.17 31.87 32.41 42.69 29.23
llama2-7B 16.08 17.67 18.47 13.88 13.55 19.46 19.44 13.02
llama2-13B 27.13 29.80 32.07 22.90 23.88 27.24 32.72 22.55
@@
Table 1 summarizes the latency results for different models and methods on the MT-Bench and Alpaca datasets. ADED consistently demonstrates lower latency compared to other methods, particularly for the vicuna-7B and llama2-13B models. For instance, on the MT-Bench dataset, ADED achieves a latency of 12.95 ms for vicuna-7B, which is significantly lower than REST (16.05 ms), REST Single Thread (16.91 ms), and Lookahead (20.53 ms). It is very important to note that the memory required for ADED (574MB) is only 5.6% of that required for REST (12GB). This trend is also observed on the Alpaca dataset, where ADED achieves a latency of 13.18 ms for vicuna-7B, compared to 13.21 ms for REST, 13.51 ms for REST Single Thread, and 20.86 ms for Lookahead.
(a) Accept Length on MT-Bench
(b) Accept Length on Alpaca
(c) Speed Up on MT-Bench
(d) Speed Up on Alpaca
Figure 3: Comparison of Accept Length and Speed Up for Different Models on (a) MT-Bench and (b) Alpaca. The performance of ADED is evaluated against REST, REST Single Thread, and Lookahead across all models and benchmarks. ADED consistently shows improvements in both accept length and speed up, demonstrating its effectiveness and robustness in different scenarios.
@@
The accept length results indicate the quality of the generated outputs, with longer accept lengths suggesting more coherent and contextually relevant text. Our method, ADED, outperforms other methods across different models on both MT-Bench and Alpaca datasets. For example, in MT-Bench, ADED achieves the highest accept length for vicuna-33B and llama2-13B models, showcasing its superior language generation capabilities.
@@
Additionally, speed up metrics are evaluated to determine the efficiencly of each method. ADED consistently shows a significant improvement in speed up across all models and datasets. This efficiency is particularly noticeable in the MT-Bench and Alpaca datasets, where ADED not only reduces latency but also enhances the overall processing speed, thus validating the robustness of our retrieval algorithms. For instance, ADED achieves a speed up of 2.4x on the MT-Bench dataset with the vicuna-13B model, outperforming REST, REST Single Thread, and Lookahead.
Figure 4: Comparison of speed up and ADED shows notable improvements.
Further analysis on the HumanEval dataset reveals that ADED achieves considerable speed up improvements. The speed up results highlight the efficiency of ADED in handling larger datasets without compromising on retrieval time or output quality. The vicuna-13B model, for example, demonstrates a speed up of nearly 2.5x when using ADED, which is a substantial improvement over the baseline methods.
These results collectively indicate that ADED provides a balanced improvement in both the quality of generated outputs and the efficiency of the retrieval process. The ability of ADED to maintain high performance across diverse tasks and datasets underscores its versatility and reliability for real-world applications. The improvements in latency, accept length, and speed up metrics affirm the efficacy of ADED in delivering superior performance while managing larger datasets effectively.
4.3 Stability of ADED
@@
In this section, we analyze the stability of our algorithm, ADED, across different categories of tasks. The categories considered include writing, roleplay, reasoning, math, coding, extraction, STEM, and humanities. The experimental results, as shown in Table 2, indicate that ADED maintains consistent performance across all categories. The average accept length remains stable, demonstrating that ADED can effectively handle a diverse range of tasks without significant variations in performance.
To further evaluate the robustness of ADED, we examined the effects of varying the top-p and temperature parameters on its performance. Figures 5(a) and 5(b) illustrate the impact of these parameters on the average accept length.
Table 2: Average Accept Length for the Tasks.
Task Avg Accept
Length
Writing 2.37
Roleplay 2.21
Reasoning 2.10
Math 2.33
Coding 2.45
Extraction 2.11
STEM 2.19
Humanities 2.17
@@
Figure 5(a) shows that changes in the top-p parameter do not significantly affect the performance of ADED. The average accept length remains relatively stable across different values of top-p, indicating that ADED is not overly sensitive to this parameter.
Similarly, Figure 5(b) demonstrates that variations in the temperature parameter have minimal impact on the performance of ADED. The consistency in the average accept length across different temperature values further supports the robustness of our algorithm.
@@
Figure 5: Sensitivity of ADED to (left) top-p and (right) temperature parameters. ADED maintains the sable acceleration under various settings.
These results confirm that ADED exhibits robust performance across a variety of tasks and maintains stability despite changes in key parameters, making it a versatile and reliable choice for diverse applications.
5 Ablation Study
To gain a deeper understanding of our method, we conduct a series of ablation studies and analyses focused on each individual component. Please see full ablation studies in the Appendix.
Figure 6: Adaptive Strategy Comparison on MTBench: Performance comparison of Vicuna-7B model with and without the adaptive strategy on the MT-Bench dataset, showing the advantage of using the adaptive approach.
Effect of the adaptive strategy. Figure 6 illustrates the performance impact of our adaptive strategy on two models, Vicuna-7B and Vicuna-13B, with a comparative analysis of average accepted lengths over varying token counts. The graphs show that the adaptive strategy consistently maintains higher average accepted lengths across the input range for both models, compared to the non-adaptive. The adaptive strategy’s success can be attributed to its dynamic adjustment of the model’s probability distributions based on the tri-gram frequencies from prior outputs. This allows the model to better manage longer contexts and maintain relevance, enhancing stability and coherence in longer interactions. The marked performance improvement, particularly in managing larger token counts, highlights the adaptive strategy’s efficacy in sustaining effective and coherent outputs in extended sequences.
Effect of the corpus size.
Table  3 demonstrates how increasing the corpus size from 300k tokens to 700k tokens affects various performance metrics. With the expansion of the corpus, there is a gradual improvement in the ’Accept Length’ from 2.17 to 2.33. This increase suggests that larger datasets provide a broader array of language patterns, which enhances the model’s ability to generate more coherent and contextually relevant outputs. Despite the growth in data size, from 253 MB to 574 MB, and a slight increase in retrieval time from 2.1 ms to 3.1 ms, the system maintains efficient data handling.
@@ 5 Ablation Study
Corpus corpus Retrieval Accept Speed
Size Size Time Length Up
300k 253 MB 2.1 ms 2.17 1.93
500k 467 MB 2.7 ms 2.24 2.01
700k 574 MB 3.1 ms 2.33 2.07
@@
The modest rise in retrieval time underscores the efficiency of the retrieval algorithms, which manage larger datasets without significantly compromising response speed. Overall, the results indicate that larger corpus sizes improve the model’s output quality while maintaining good system performance.
These findings highlight the importance of the size of the corpus in enhancing the performance of ADED, demonstrating that a larger and more carefully selected corpus can significantly improve the quality and efficiency of the generated outputs.
6 Related Work
Approach Without Draft Models. A significant portion of recent advances in language model decoding strategies has focused on improving efficiency without relying on draft models. Two notable approaches in this realm are Lookahead decoding Fu et al. (2024a) and Retrieval-Based Speculative Decoding He et al. (2023). Lookahead decoding is an approach that enhances the efficiency of the decoding process through the prediction of subsequent tokens via Jacobi Iteration Sleijpen and Van der Vorst (2000). It employs a heuristic to estimate the future cost of a sequence without the need to explicitly create a draft. This technique not only accelerates the decoding process by minimizing the number of tokens to be processed but also seeks to preserve or improve the quality of the output text by taking into account potential future scenarios in the decision-making process. Retrieval-Based Speculative Decoding(REST) He et al. (2023) introduces a retrieval-enhanced generation model that speculatively decodes sequences without the need for producing preliminary drafts. It instead searches and prioritizes possible continuations from an already established sequence database. This approach utilizes the inherent redundancy in natural language to anticipate probable continuations, thereby eliminating the requirement for draft sequences and greatly reducing computational costs. Both Lookahead decoding and REST demonstrate the capabilities of decoding methods that avoid intermediate drafts, providing a more straightforward and computationally efficient route to generating high-quality text.
@@ 6 Related Work
Approach With Draft Models. Draft models are also used to improve the decoding efficiency. Techniques such as Speculative Decoding (Leviathan et al., 2023; Spector and Re, 2023; Chen et al., 2023; Stern et al., 2018), Madusa Cai et al. (2024), Eagle Li et al. (2024b), various other approaches requiring draft models (Zhang et al., 2024; Liu et al., 2023b; Kim et al., 2024; Fu et al., 2024b) fall into this category, utilizing models to generate drafts. Although these methods aim to speed up response times and reduce computational load during initial text generation, their adoption comes with significant drawbacks. The primary issue is the necessity for additional training specific to the draft models, which can be resource-intensive. Moreover, these techniques generally depend on GPU resources (Kwon et al., 2023; Sheng et al., 2023; Park et al., 2024) for inference, potentially limiting their application in environments where such hardware is unavailable or when operating under strict resource constraints. This dependence not only increases operational costs, but also restricts flexibility in deployment scenarios.
7 Conclusion

### v2

@@ Introduction
where $P$ denotes the conditional probability of LLM, $w$ denotes the tokens, and $\theta$ denotes the model parameters.
@@
Table 1: Latency and Average Accept Length Comparison between ADED and Baselines. In most test cases, ADED has the lowest latency, longer accept length, and higher efficiency.
Latency Average Accept Length
Benchmark Model REST REST Single Lookahead Autoregressive ADED REST REST Single Lookahead ADED
MT-Bench Vicuna-7B 16.31 17.36 18.93 24.77 12.95 1.97 1.98 1.89 2.42
Vicuna-13B 25.43 25.99 32.73 44.07 22.94 1.98 1.99 1.85 2.39
Vicuna-33B 28.63 28.62 40.53 52.97 24.96 1.95 1.96 1.83 2.29
Llama2-7B 16.08 17.67 18.84 25.58 13.85 1.96 1.95 1.96 2.30
Llama2-13B 27.13 29.80 31.24 44.76 25.13 1.95 1.95 1.96 2.32
Alpaca Vicuna-7B 14.24 14.58 18.73 24.49 12.81 2.22 2.22 1.89 2.33
Vicuna-13B 22.94 23.01 32.60 43.60 24.06 2.21 2.21 1.86 2.26
Vicuna-33B 26.03 25.89 40.58 52.52 24.62 2.11 2.12 1.82 2.21
Llama2-7B 14.13 14.87 19.28 25.38 12.90 2.21 2.20 1.97 2.37
Llama2-13B 23.66 24.07 31.18 44.04 23.57 2.15 2.13 1.96 2.32
Human Eval Vicuna-7B 14.90 15.56 18.99 25.49 11.24 2.21 2.23 2.10 2.67
Vicuna-13B 20.17 20.61 27.43 45.13 19.96 2.50 2.50 2.23 2.81
Vicuna-33B 24.91 25.06 31.34 52.32 21.19 2.29 2.30 2.02 2.62
Llama2-7B 14.37 15.57 15.28 25.91 11.68 2.19 2.19 2.27 2.63
Llama2-13B 25.46 25.85 26.72 45.25 21.82 2.01 2.01 2.17 2.60
(a) Throughput and Speedup on MT-Bench
(b) Throughput and Speedup on Alpaca
(c) Throughput and Speedup on Human-Eval
Figure 3: Comparison of ADED’s throughput for different models on (a) MT-Bench, (b) Alpaca, and (c) Human-Eval. The performance of ADED shows stable and significant improvements across different models and benchmarks.
Comparative Analysis: As shown in Eq (6) and Eq. (7), both MCTS and LLMs can be viewed as regularized optimization problems for selecting the distribution of the next tokens. On the one hand, the Q-function in MCTS for ADED can be viewed as an approximation to the log-likelihood of LLMs:
$Q(x,a)$ $=-\sum_{t=2}^{T}\log\hat{P}(w_{t}\mid w_{t-1},w_{t-2};\theta)$
$\approx\log P(w_{0},w_{1},\cdots,w_{T};\theta)$
$=-\sum_{t=2}^{T}\log P(w_{t}\mid w_{1:t-1};\theta),$ (8)
where $\hat{P}$ and $P$ are the conditional probability distribution from tri-gram-matrix-based LLM representative and LLMs, respectively. On the other hand, both MCTS and LLMs employ regularization to improve the optimization procedure. As a result, we verify the similarities between MCTS and LLM Inference in terms of optimization and regularization.
Experiments
Experimental Setup
Models and Datasets. We conduct a series of experiments with five distinct models on three datasets to evaluate the efficacy of ADED. In particular, We use three Vicuna models (Chiang et al. 2023) (7B, 13B, 33B) and two LLaMA2-chat models (Touvron et al. 2023b) (7B, 13B) to evaluate the acceleration capabilities across different model sizes and types. Our assessment incorporates the HumanEval (Chen et al. 2021), MT-Bench (Zheng et al. 2023), and Alpaca (Taori et al. 2023) datasets to ascertain general natural language understanding and generation competencies. These datasets are meticulously chosen to guarantee a comprehensive analysis of the acceleration techniques across various tasks.
Corpus. We construct two corpora. The first one is built using a portion of the Python pre-training code from The Stack (Kocetkov et al. 2022), comprising about 2.7M Python code samples with a resulting size of 1007MB. The second is constructed using data derived from UltraChat (Ding et al. 2023), consisting of around 774K ChatGPT conversations, producing a corpus with a size of 574MB. The experiments on the MT-Bench and Alpaca are conducted using the UltraChat corpus, while the Human-Eval benchmark utilize the corpus from The Stack.
Metrics. To assess the acceleration performance on large language models, we use two main metrics: speedup ratio and average acceptance length. Speedup ratio, calculated as the ratio of the time required by the baseline models to complete inference tasks without acceleration to the time required by our ADED, measures the efficiency gains introduced by the algorithm. The second metric, average acceptance length, measures the average number of tokens accepted per forward pass by the target large language models, excluding any overhead of retrieving and constructing draft tokens, indicating the maximum possible acceleration.
Baselines. We compare various foundational approaches to improve the decoding speed of large language models. We examine Lookahead Decoding (Fu et al. 2024a), a precise and parallel decoding algorithm that cuts down latency without relying on draft models. We compare REST (He et al. 2024) (Retrieval-Based Speculative Decoding), which adopts a retrieval-based strategy to create draft tokens, in contrast to conventional speculative decoding methods that rely on a draft model.For fairness in comparison, we include REST Single, a single-threaded version of REST, to evaluate performance under constrained processing conditions. We also include the traditional Autoregressive method, which represents the standard decoding approach, serving as a baseline to highlight the improvements offered by the other methods. All experiments are conducted on NVIDIA A6000 GPUs, except for the 33B model, which utilizes an NVIDIA H100. The experiments default to Greedy sampling.
Configuration of ADED. To ensure reproducibility, we provide detailed hyperparameters, the corpus used, and the runtime environment for each set of experiments in Appendix F. This allows interested researchers to replicate our findings accurately and compare them against their setups. Due to space constraints, these details are not included in the main text but are comprehensively documented in the appendix.
Main Results
@@
In the experiments, we compare the efficacy of different baselines applied to various models, utilizing three datasets: MT-Bench, Human-Eval, and Alpaca. We focus on metrics of Accept Length, Latency, and Speedup Ratio. Table 1 summarizes the latency and average accept length on the three datasets. ADED consistently demonstrates lower latency, particularly for the vicuna-7B and llama2-13B models. For instance, on MT-Bench, ADED achieves a latency of 12.95 ms for vicuna-7B, which is lower than REST (16.31 ms), REST Single Thread (17.36 ms), and Lookahead (18.93 ms). Notably, the memory required for ADED (574MB) is only 5.6% of that required for REST (12GB). According to Table 2, even when ADED uses a smaller corpus (253MB, which is just 2.5% of REST’s requirements), it still achieves lower latency than REST. This trend is also observed on Alpaca, where ADED achieves a latency of 12.81 ms for vicuna-7B, compared to 14.24 ms for REST, 14.58 ms for REST Single Thread, and 18.73 ms for Lookahead.
@@
The accept length results in Table 1 indicate the quality of the generated outputs, with longer accept lengths suggesting more coherent and contextually relevant text. Our method, ADED, outperforms other methods across different models on both MT-Bench and Alpaca datasets. For example, on MT-Bench, ADED achieves the highest accept length for vicuna-33B and llama2-13B models, showcasing its superior language generation capabilities.
@@
Speedup ratio are used to evaluate the efficiency. ADED consistently shows a significant improvement in speed up across all datasets in Figure  3. This efficiency is noticeable on the MT-Bench, Alpaca and Human-Eval datasets, where ADED not only reduces latency but also enhances the overall processing speed. For instance, ADED achieves a speedup of 1.92x on MT-Bench with the vicuna-13B model, outperforming REST, REST Single Thread, and Lookahead. On the HumanEval dataset, the vicuna-33B model, for example, demonstrates a speedup of nearly 2.5x when using ADED.
Stability of ADED
(a) Effect of the adaptive strategy
(b) Stability amount different tasks
Figure 4: (4(a)) Adaptive Strategy comparison on MTBench: Performance of Vicuna-7B model with and without the adaptive strategy on the MT-Bench dataset, showing the advantage of using the adaptive approach. (4(b)) Average Accept Length for different tasks on MT-Bench, demonstrating that ADED consistently performs well across tasks.
@@
In this section, we analyze the stability of ADED across different categories of tasks. The categories include writing, roleplay, reasoning, math, coding, extraction, STEM, and humanities. The experimental results shown in Figure 4(b) indicate that ADED maintains consistent performance across categories. The average accept length remains stable, demonstrating that ADED can effectively handle a diverse range of tasks without significant variations in performance.
@@
To further evaluate the robustness of ADED, we examine the effects of varying the top-p and temperature parameters on the performance. Figures 5(a) and 5(b) summarize the impact of these parameters on the average accept length. In particular, Figure 5(a) shows that changes of top-p do not affect the performance of ADED significantly. The average accept length remains relatively stable across different values of top-p, indicating that ADED is not overly sensitive to this parameter. Figure 5(b) demonstrates that variations in the temperature parameter have negligible impact on the performance of ADED. The consistency in the average accept length across different temperature values further supports the robustness.
@@
Figure 5: Sensitivity analysis of ADED on (5(a)) top-p and (5(b)) temperature parameters.
Ablation Study
To gain insight into our method, we conduct a series of ablation studies. Full studies are summarized in Appendix D.
Effect of the adaptive strategy. Figure 4(a) illustrates the performance of our adaptive strategy on Vicuna-7B, with the analysis of average accept lengths over varying token counts. We find that the adaptive strategy maintains a higher average accept length over the entire range compared to the non-adaptive strategy. The adaptive strategy’s success is attributed to its dynamic adjustment of the model’s probability distributions based on the tri-gram frequencies from prior outputs. This allows the model to better manage longer contexts and maintain relevance, enhancing stability and coherence in longer interactions.
Effect of the corpus size. Table 2 shows the impact of the increase in corpus size from 121k to 774k on various performance metrics. With the expansion of the corpus, there is a gradual improvement in the Accept Length from 2.30 to 2.42. This increase suggests that larger datasets provide a broader array of language patterns, which enhances the model’s ability to generate more coherent and contextually relevant outputs. Despite the increase in data size from 253 MB to 574 MB, the system maintains efficient data processing capability. The small differences in latency affirm ADED’s consistent performance, even with smaller corpus sizes, which further extends its potential for use on resource-constrained devices.
@@
Corpus Size Corpus Size Latency Accept Length Speed up
121k 253 MB 13.09 ms 2.30 1.89
774k 574 MB 12.95 ms 2.42 1.93
@@
The modest rise in retrieval time underscores the efficiency of the retrieval algorithms, which can manage larger datasets without significantly compromising response speed. In summary, the results show that larger corpus sizes can improve the quality of the model’s output while maintaining good system performance.
Effect of MCTS. Figure 6 presents the results for Vicuna-7B and Vicuna-13B models on the MT-Bench dataset, showing the impact of different MCTS search counts on performance. We find that for both models, increasing the number of searches improves performance, while the optimal number varies by model size. The average exact length and latency are plotted against the number of searches, illustrating the trade-off between performance and computational cost. As shown in the figure, there is a notable improvement in the average exact length as the number of searches increases, along with an increase in latency, indicating a balance between search depth and generation time. We further compare greedy search with MCTS (full traversal is not feasible due to the huge search space) while keeping the number of 150 search iterations constant. The results show that the average accept length of the greedy search is only 1.493, significantly lower than that obtained by MCTS, demonstrating the superiority of MCTS in efficiently managing the vast decision spaces.
(a) Vicuna-7B
(b) Vicuna-13B
Figure 6: Comparison of Search Counts and Performance on MT-Bench.
Effect of N-gram Model Choice. Our studies extensively evaluate the impact of different n-gram configurations on decoding performance. In tests conducted on the MT-Bench dataset using the Vicuna-7B model, bi-grams and 4-grams result in accept lengths of 1.80 and 1.82, respectively. These results are significantly lower compared to the 2.30 accept length achieved with tri-grams. Bi-gram models demonstrate limited capability in effectively utilizing contextual information, often leading to outputs that appear more random and less coherent. Conversely, 4-grams exhibit overly deterministic behavior, constraining the diversity of the generated text due to their restrictive nature in capturing extensive prior context. Tri-grams strike an optimal balance, providing enough contextual depth to enhance the coherence and relevance of outputs while still allowing for sufficient variability and diversity in text generation. This balance makes tri-grams particularly effective in large language model decoding, as they encapsulate the ideal compromise between randomness and contextual awareness.
Related Work
@@
A number of research efforts on decoding strategies for large language models have used draft models to improve decoding efficiency. Techniques such as Speculative Decoding (Leviathan, Kalman, and Matias 2023; Spector and Re 2023; Chen et al. 2023; Stern, Shazeer, and Uszkoreit 2018), Madusa (Cai et al. 2024), Eagle (Li et al. 2024b), various other approaches requiring draft models (Zhang et al. 2024; Liu et al. 2023b; Kim et al. 2024; Fu et al. 2024b) fall into this category, utilizing models to generate drafts. Specifically, Speculative Decoding uses an advanced sampling technique where the auxiliary model generates a set of potential token sequences, and the primary model selects the most sequences, resulting in a good balance between speed and accuracy. Although these methods primarily aim to enhance the accuracy of generated texts and significantly accelerate the response time during initial text generation, their adoption comes with drawbacks. The primary issue is the necessity for additional training specific to the draft models, which could be resource-intensive. Moreover, these techniques generally depend on GPU resources (Kwon et al. 2023; Sheng et al. 2023; Park et al. 2024) for inference, potentially limiting their application in environments where such hardware is unavailable or when operating under strict resource constraints.
A significant portion of recent advances has focused on improving efficiency without relying on draft models (Fu et al. 2024a; He et al. 2024). Two notable approaches in this realm are Lookahead decoding (Fu et al. 2024a) and Retrieval-Based Speculative Decoding (REST) (He et al. 2024). Lookahead decoding is an approach that enhances the efficiency of the decoding process through the prediction of subsequent tokens via Jacobi Iteration (Sleijpen and Van der Vorst 2000). It employs a heuristic to estimate the future cost of a sequence without the need to explicitly create a draft. REST introduces a retrieval-enhanced generation model that speculatively decodes sequences without the need for producing preliminary drafts. It instead searches and prioritizes possible continuations from an already established sequence database. However, these methods exhibit lower accuracy and greater resource use compared to our approach. They demand more memory and GPU processing power, posing challenges in resource-scarce settings.
Conclusion